# Transfer-based Adversarial Poisoning Attacks for Online (MIMO-)Deep Receviers

### v2

@@ I Introduction
In recent years, the application of deep learning (DL) in designing wireless communication systems has garnered significant interest. Researchers have concentrated on employing DL in wireless receivers to enhance communication performance in complex channels and to bolster adaptability in dynamic environments [2, 3, 4]. However, DL-based wireless applications face vulnerabilities to evasion and data poisoning attacks owing to the inherent openness of wireless channels and the fragility of neural models [5]. Investigating attack methodologies on deep receivers serves to elucidate their response under such threats, thereby facilitating the development of secure wireless DL systems, which forms the primary focus of this paper.
@@ I-A DL-based Wireless Receivers and Related Applications
Till now, numerous studies have explored DL-based designs for wireless communication systems. Most of them utilize an independent DNN to map the input-output relationships of functional modules in communication links. Jointly optimizes modules at both the transmitter and receiver by multiple DNNs. Typically, the applications include DL-based adaptive modulation[6], channel estimation[7], channel coding and decoding[8, 9], and modulation recognition[10]. In addition to replacing functional modules in the physical layer, various constraints can be integrated into the training process to optimize additional system metrics, such as the adjacent channel leakage ratio (ACLR) and peak-to-average power ratio (PAPR) [6],[11]. For the designs of the receiver, lots of work has contributed to enhancing its adaptability to the dynamic channels, including training and integrating multiple models with extensive data [12] and joint learning [13, 14, 15].
However, DL-based wireless designs mentioned above are data-driven methods that heavily depend on a substantial amount of training data to improve their generalization capacity. Given that deep receivers typically have access to only limited pilots for adaptation, this characteristic poses a significant challenge. In addition, data-driven designs may suffer from performance degradation when faced with data distribution drift caused by dynamic channels[16]. To handle these issues, an online learning based approach has been proposed, involving dataset, training algorithm and deep receivers architecture. In particular, data augmentation[17] and self supervision methods [18],[19] were proposed to to expand the training data for online adaptation. In [16] and [20], meta learning was employed to improve the generalization capability. Furthermore, Model-based deep learning provide a solution for receivers architecture design to satisfy the adaptability and data efficiency[21]. Specifically, the deep receivers were explicitly modelled by incorporating wireless domain knowledge, thereby reducing the dependence on data, such as DNN-aided inference[15],[21] and deep unfolding [21, 15, 22, 1]. In these studies, [1] proposed a classic model-based deep receiver, i.e., the DeepSIC, in MIMO scenarios, derived from the iterative soft interference cancellation (SIC)[23] MIMO detection algorithm. It employed DNN instead of each round of interference cancellation and soft detection, requiring only a few iterations to achieve extremely low data dependence and optimal performance. [20] utilized meta-learning to improve the training performance of online DeepSIC, and the evaluation results indicated that its performance was improved compared with traditional data-driven receiver, and it exhibited commendable adaptability to dynamic channels.
@@ I-B Security of DL in Wireless Communications
As mentioned earlier, while DL-based transceiver designs can enhance performance, they remain vulnerable to attacks by malicious users. In particular, attacks on DL-based transceivers are divided into two main categories, i.e., the evasion attacks and data poisoning attacks. Evasion attacks, also known as adversarial attacks, manipulates test data to mislead the model[5],[24]. On the other hand, data poisoning attacks corrupt the training data, affecting the model’s performance during testing[5], [24, 25, 26].
According to extensive literature review, numerous studies on DL-based wireless communication primarily concentrate on evasion attacks. For instance, [27] proposed an adversarial attack method against adaptive modulation. [28] proposed generative adversarial network (GAN)-based method to generate adversarial perturbations for channel received data, which can unnoticeably mislead wireless end-to-end autoencoders, modulation pattern recognition, and the DL-based symbol detection in orthogonal frequency division multiplexing (OFDM) systems. [29] reported adversarial perturbations can interfere with gradient-based iterative optimization algorithms in the physical layer. [30] proposed semantic attacks against semantic communication. Furthermore, adversarial perturbations can also play a role in interpretable (e.g., deep unfolding-based) architectures. To illustrate, [31] employed transfer-based methods to attack deep sparse coding networks and demonstrated that these attacks exert deleterious effects on the various components of deep unfolded-based sparse coding. Regarding data poisoning attacks, current research primarily focuses on cognitive radio spectrum-aware poisoning [5],[32] and disrupting distributed wireless federated learning[33, 34].
@@ I-C Contribution of This Paper
Unlike previous studies, this paper addresses security threats to online deep receivers. Furthermore, we propose a transfer-based adversarial poisoning attack method, which can significantly corrupt various online deep receivers even without prior knowledge of the target system. Specifically, we focus on online receivers based on model-based deep learning, such as DeepSIC[1] and Meta-DeepSIC[20], as well as general DNN detectors, including the black-box DNN detector[15],[17] and the ResNet detector designed based on DeepRX[35].
As previously stated, DeepSIC is a classical model-based deep receiver that can be combined with meta learning for efficient online adaptation. This design effectively tackles the challenge of limited pilot data in wireless communication scenarios, thereby improving the generalization of deep receivers under dynamic channel conditions. Moreover, studies on the attack methods for DeepSIC can provide comprehensive insights into deep receiver characteristics and contribute to robust designs. Ultimately, this research aids in creating secure and efficient DL-enabled wireless communication systems.
@@
    We highlight a communication system susceptible to malicious user poisoning attacks. We then analyze the vulnerability of the deep receiver based on online learning in the authorization system. From the perspective of malicious users, we further develop an attack utility model and an optimal attack utility decision problem.
@@
    We effectively design a poisoning attack framework and attack perturbation generation method for online learning deep receivers. The fundamental concept is to introduce a poisoned sample into the online training and updating phase of the deep receiver, thereby compromising its performance over time. The poisoning attack framework has two stages. Firstly, malicious users employ joint learning to create a surrogate model, which can be selected from a generic DNN architecture, e.g., feedforward DNNs. Secondly, they generate poisoning perturbation samples based on the surrogate model. The transferability of the poisoning attack makes it work on different types of deep receivers.
@@ II-A Communication System Scenario Model
In this paper, we investigate a poisoning attack scenario model for a communication system, as illustrated in Fig. 1. This system consists of a pair of legitimate transceivers and a malicious poisoning attack user. Both the legitimate transmitter and receiver equip with multiple antennas, denoted as $N_{tx}$ and $N_{rx}$ respectively. We focus on a single-antenna malicious user, as this represents a cost-effective and straightforward approach to conducting attacks. The data transmission from transceiver to receiver is block-based, as illustrated in Fig. 2. The length of one block is $L$, including $L_{\text{pilot}}$ pilot symbols and $L_{\text{info}}$ information symbols, herein, $L_{\text{info}}\gg L_{\text{pilot}}$. As shown in Fig. 3, the legitimate receiver utilizes a DL-based architecture for signal receiving and processing. It is trained using pilot data and employs the trained deep receiver to decode information data. For the malicious user, based on the previously collected pilot data, it launches an attack by poisoning or disturbing the transmission process of the pilot used by the legitimate user. Its objective is to corrupt with the online training and updating of the deep receiver, thereby disrupting its information data reception and decoding.
@@ 4. Step 4:
In principle, a poisoning attack perpetrated by a malicious user can be conceptualized as a perturbation injection process. In particular, the perturbation is defined as a vector in the complex space of receiver inputs, denoted as $\delta\in\mathbb{C}^{N_{rx}}$, with a poisoning process represented by $\mathcal{P}(\cdot)$. For the $i$-th received symbol $\mathbf{y}_{i}$, the corresponding poisoning perturbation is given by $\delta_{i}$, and the corresponding poisoned received symbol is given by $\mathcal{P}\left(\mathbf{y}_{i}\right)=\mathbf{y}_{i}+\delta_{i}$.
Therefore, within the context of poisoning attacks on deep receivers, the primary challenge for malicious users is to design an optimal perturbation signal structure that maximizes the deep receiver’s loss on subsequent information symbols or validation sets, which will be addressed in the following sections.
@@ III Adversarial Machine Learning Theory
Before introducing the poisoning attack method for online deep receivers proposed in this paper, we briefly discuss the theory of adversarial machine learning. In particular, we provide a brief overview of the threat model presented in Section III-A, including the concepts of the attacker’s goal, the attacker’s knowledge, and the attacker’s capability. The aforementioned concepts and definitions facilitate a more comprehensive understanding of the attack method for deep receivers proposed in this paper. Subsequently, in Section III-B and Section III-C, the fundamental optimization issues associated with two distinct attack paradigms, namely evasion attacks and data poisoning attacks, are briefly elucidated. Furthermore, we explain the differences and connections between the two attack paradigms. Finally, we analyze the transferability of attack samples in Section III-D, and discuss the methods to enhance the transferability of attacks.
III-A Threat Model: Attacker’s Goal, Knowledge and Capabilities
@@ III-A1 Attacker’s Goal
As discussed in [24],[36],[37], the objective of the attacker in adversarial machine learning scenarios can be categorised according to the form of security threats, including integrity attacks and availability attacks.
@@
    Integrity attacks: The attacker’s goal is to tamper the integrity of the target. Specifically, this implies that the attack samples generated by the attacker are only effective in certain parts of the target system, while the remainder of the target system retains its original functionality. For example, an attacker applies a specific adversarial perturbation to an image of a traffic sign in order to circumvent the detection of the visual classifier in an automated vehicle. However, at that moment, the visual classifier remains accurately operational for other objects.
@@
    Availability attacks: In contrast, availability attacks aim to disrupt the normal functioning of the entire system, rendering it unavailable to legitimate users.
The difference between integrity attacks and availability attacks primarily arises from the different focuses of the optimization objectives inherent to the attack models constructed. In this paper, the proposed attack method is an availability attack, namely the destruction of the usability of all functions of a deep receiver.
@@ III-A2 Attacker’s Knowledge
The knowledge possessed by the attacker is indicative of the extent to which they are aware of the attack target. This knowledge encompasses several key dimensions, as outlined in [24],[36]. These dimensions include: (i) The data utilized for training purposes. (ii) The architectural design of the target model, the learning algorithms employed during training, along with their associated parameters and training hyperparameters. (iii) The data comprising the test set. Based on the combination of these dimensions, two main attack scenarios can be further defined as below:
@@
    White-box attacks: We consider that the attacker has complete knowledge about the attack target. In this context, the attacker will adapt the nature of the attack to align with the specific characteristics of the target to achieve the most effective and impactful outcome.
@@
    Black-box attacks: Black box attacks can be further categorized into two main types: Transfer-based attacks and query attacks. In transfer-based attacks, the attacker lacks or only has partial knowledge of the internal workings of the target model. The knowledge about the target encompasses the aforementioned dimensions (i), (ii), and (iii). In this setting, the attacker is limited to relying on the data they have collected to construct a surrogate model that approximates the target model. This attack is then transferred to the target model by launching a white-box attack on the surrogate model. In a black-box query attack, the attacker can query the target’s output or confidence level to optimize the attack. Currently, the majority of black-box attacks exploit the transferability for attack purposes [24],[38, 39, 40, 41].
@@ III-A3 Attacker’s Capability
The attacker’s capabilities determine the methods used to influence the attack target and the specific constraints for data manipulation. To avoid potential defense filtering mechanisms, the attacker must impose constraints on the manipulated data. It is standard practice to add an upper bound of $\epsilon$ to the perturbation $\delta$ under the $p$-norm space. From the perspective of attack methods[24], if the attacker can manipulate data from both the training and testing phases simultaneously, which considered as causal attacks and is called data poisoning attacks. If the attacker can only manipulate the data during the testing phase, this attack is considered exploratory and is called evasion attacks. Their difference lies in solving optimization objectives and implementation methods. This paper focuses on adversarial poisoning attacks in black-box scenarios, which can be seen as a synthesis of evasion attacks and data poisoning attacks. The specific optimization goals and implementation forms of evasion attacks and data poisoning attacks are described in the following Section III-B and Section III-C, respectively.
@@ III-B Evasion Attacks
The evasion attacks are implemented by constructing adversarial samples whose objective is to induce a catastrophic misclassification in the testing phase of the DNN. This is achieved by identifying the blind spots of the DNN and introducing carefully crafted tiny perturbations to the input, as illustrated in Fig. 5(a). The construction of adversarial samples arises from the observation that the deeper features and outputs of a classifier can undergo notable alterations when the inputs undergo slight directional changes. Gradient-based optimizers [42],[43] are capable of readily identifying these directions, which exert an influence on the DNN. Accordingly, the methodology for the construction of adversarial samples can be expressed as follows: In the case of a given target model and inputs, the gradient of the objective function is employed to direct the application of minor perturbations to the input data to maximise the loss incurred by the input. This process can therefore be conceptualised as a single-layer optimization problem. Specifically, $\mathbf{y}$ denotes the model input, $\mathbf{s}$ denotes the labels corresponding to $\mathbf{y}$, and $\mathbf{y}^{\prime}$ denotes the adversarial samples after the adversarial perturbation is added. The adversarial perturbation $\delta=\left\|\mathbf{y}^{\prime}-\mathbf{y}\right\|_{p}$ has an upper bound of $\epsilon>0$, and the optimal adversarial sample obtained after optimization is $\mathbf{y}^{*}$. $\theta$ denotes the parameters of the classifier $f_{\theta}$. For the classification task, the cross-entropy loss function $\mathcal{L}\left(\mathbf{y}^{\prime},\mathbf{s};f_{\theta}\right)$ is defined. Thus, finding the optimal adversarial samples constitutes the following optimization problem

### v3

@@ I Introduction
In recent years, the application of deep learning (DL) in designing wireless communication systems has garnered significant interest. Researchers have concentrated on employing DL in wireless receivers to enhance communication performance in complex channels and to bolster adaptability in dynamic environments [3, 4]. However, DL-based wireless applications face vulnerabilities to evasion and data poisoning attacks owing to the inherent openness of wireless channels and the fragility of neural models [5]. Investigating attack methodologies on deep receivers serves to elucidate their response under such threats, thereby facilitating the development of secure wireless DL systems, which forms the primary focus of this paper.
@@ I-A DL-based Wireless Receivers and Related Applications
Till now, numerous studies have explored DL-based designs for wireless communication systems. Most of them utilize an independent DNN to map the input-output relationships of functional modules in communication links. Alternatively, some approaches jointly optimize modules at both the transmitter and receiver by multiple DNNs. Typical applications include DL-based adaptive modulation[6], channel estimation[7], channel coding and decoding[8], and modulation recognition[10]. Beyond replacing functional modules in the physical layer, various constraints can be integrated into the training process to optimize additional system metrics, such as the adjacent channel leakage ratio (ACLR) and peak-to-average power ratio (PAPR) [6]. In terms of receiver design, lots of work has contributed to enhancing its adaptability to the dynamic channels, including training multiple models as a deep ensemble [12] and joint learning [13, 15].
However, DL-based wireless designs mentioned above are data-driven methods that heavily depend on a substantial amount of training data to improve their generalization capacity. Given that deep receivers typically have access to only limited pilots for adaptation, this characteristic poses a significant challenge. In addition, data-driven designs may suffer from performance degradation when faced with data distribution drift caused by dynamic channels[16]. To handle these issues, an online learning based approach has been proposed, involving dataset, training algorithm and deep receivers architecture. In particular, data augmentation[17] and self supervision methods [18],[19] were proposed to to expand the training data for online adaptation. In [16] and [20], meta-learning was employed to improve the generalization capability. Furthermore, Model-based deep learning provide a solution for receivers architecture design to satisfy the adaptability and data efficiency[21]. Specifically, the deep receivers were explicitly modelled by incorporating wireless domain knowledge, thereby reducing the dependence on data, such as DNN-aided inference[15],[21] and deep unfolding [21, 15, 22, 1]. In these studies, [1] proposed a classic model-based deep receiver, i.e., the DeepSIC, in MIMO scenarios, derived from the iterative soft interference cancellation (SIC)[23] MIMO detection algorithm. It employed DNN instead of each round of interference cancellation and soft detection, requiring only a few iterations to achieve extremely low data dependence and optimal performance. [20] utilized meta-learning to improve the training performance of online DeepSIC, and the evaluation results indicated that its performance was improved compared with traditional data-driven receiver, and it exhibited commendable adaptability to dynamic channels.
@@ I-B Security of DL in Wireless Communications
As mentioned earlier, while DL-based transceiver designs can enhance performance, they remain vulnerable to attacks by malicious users. In particular, attacks on DL-based transceivers are divided into two main categories, i.e., the evasion attacks and data poisoning attacks. Evasion attacks, also known as adversarial attacks, manipulate test data to mislead the model[5],[24]. On the other hand, data poisoning attacks corrupt the training data, affecting the model’s performance during testing[5], [24, 25, 26].
According to extensive literature review, numerous studies on DL-based wireless communication primarily concentrate on evasion attacks. For instance, [28] proposed generative adversarial network (GAN)-based method to generate adversarial perturbations for channel received data, which can unnoticeably mislead wireless end-to-end autoencoders, modulation pattern recognition, and the DL-based symbol detection in orthogonal frequency division multiplexing (OFDM) systems. [29] reported adversarial perturbations can interfere with gradient-based iterative optimization algorithms in the physical layer. [30] proposed semantic attacks against semantic communication. Furthermore, adversarial perturbations can also play a role in interpretable (e.g., deep unfolding-based) architectures. To illustrate, [31] employed transfer-based methods to attack deep sparse coding networks and demonstrated that these attacks exert deleterious effects on the various components of deep unfolded-based sparse coding. Regarding data poisoning attacks, current research primarily focuses on cognitive radio spectrum-aware poisoning [5],[32] and disrupting distributed wireless federated learning[33, 34].
@@ I-C Contribution of This Paper
Unlike previous studies, this paper addresses security threats to online deep receivers. Furthermore, we propose a transfer-based adversarial poisoning attack method, which can significantly corrupt various online deep receivers even without prior knowledge of the target system. Specifically, we focus on online receivers based on model-based deep learning, such as DeepSIC[1] and Meta-DeepSIC[20], as well as general DL-based detectors, including the black-box DNN detector[15],[17] and the ResNet detector designed based on DeepRX[35].
As previously stated, DeepSIC is a classical model-based deep receiver that can be combined with meta-learning for efficient online adaptation. This design effectively tackles the challenge of limited pilot data in wireless communication scenarios, thereby improving the generalization of deep receivers under dynamic channel conditions. Moreover, studies on the attack methods for DeepSIC can provide comprehensive insights into deep receiver characteristics and contribute to robust designs. Ultimately, this research aids in creating secure and efficient DL-enabled wireless communication systems.
@@
    We highlight a communication system susceptible to malicious user poisoning attacks. We then analyze the vulnerability of the deep receiver based on online learning in the authorization system. From the perspective of malicious user, we further develop an attack utility model and an optimal attack utility decision problem.
@@
    We effectively design a poisoning attack framework and attack perturbation generation method for online learning deep receivers. The fundamental concept is to introduce a poisoned sample into the online training and updating phase of the deep receiver, thereby compromising its performance over time. The poisoning attack framework has two stages. Firstly, malicious user employ joint learning to create a surrogate model, which can be selected from a generic DNN architecture, e.g., feedforward DNNs. Secondly, they generate poisoning perturbation samples based on the surrogate model. The transferability of the poisoning attack makes it work on different types of deep receivers.
@@ II-A Communication System Scenario Model
In this paper, we investigate a poisoning attack scenario model for a communication system, as illustrated in Fig. 1. This system consists of a pair of legitimate transceivers and a malicious poisoning attack user. Both the legitimate transmitter and receiver equip with multiple antennas, denoted as $N_{tx}$ and $N_{rx}$ respectively. We focus on a single-antenna malicious user, as this represents a cost-effective and straightforward approach to conducting attacks. The data transmission from transmitter to receiver is block-based, as illustrated in Fig. 2. The length of one block is $L$, including $L_{\text{pilot}}$ pilot symbols and $L_{\text{info}}$ information symbols, herein, $L_{\text{info}}\gg L_{\text{pilot}}$. As shown in Fig. 3, the legitimate receiver utilizes a DL-based architecture for signal receiving and processing. It is trained using pilot data and employs the trained deep receiver to decode information data. For the malicious user, based on the previously collected pilot data, it launches an attack by poisoning or disturbing the transmission process of the pilot used by the legitimate user. Its objective is to corrupt with the online training and updating of the deep receiver, thereby disrupting its information data reception and decoding.
@@ 4. Step 4:
In principle, a poisoning attack perpetrated by a malicious user can be conceptualized as a perturbation injection process. In particular, the perturbation is defined as a vector in the complex space of receiver inputs, denoted as $\delta\in\mathbb{C}^{N_{rx}}$, with a poisoning process represented by $\mathcal{P}(\cdot)$. For the $i$-th received symbol $\mathbf{y}_{i}$, the corresponding poisoning perturbation is given by $\delta_{i}$, and the corresponding poisoned received symbol is given by $\mathcal{P}\left(\mathbf{y}_{i}\right)=\mathbf{y}_{i}+\delta_{i}$. Therefore, within the context of poisoning attacks on deep receivers, the primary challenge for malicious users is to design an optimal perturbation signal structure that maximizes the deep receiver’s loss on subsequent information symbols or validation sets, which will be addressed in the following sections.
@@ III Adversarial Machine Learning Theory
Before introducing the proposed poisoning attack method for online deep receivers, we briefly discuss the theory of adversarial machine learning. Specifically, we provide a overview of the threat model presented in Section III-A, including the attacker’s goal, knowledge, and capability. These concepts and definitions facilitate a more comprehensive understanding of the proposed attack method for deep receivers. Subsequently, in Section III-B and Section III-C, we briefly elucidate the fundamental optimization issues associated with two distinct attack paradigms, namely evasion attacks and data poisoning attacks. Furthermore, we explain the differences and connections between these two attack paradigms. Finally, in Section III-D, we analyze the transferability of attack samples and discuss the methods to enhance their transferability.
III-A Threat Model
@@ III-A1 Attacker’s Goal
As discussed in [24],[36],[37], the objective of the attacker in adversarial scenarios can be categorised according to the form of security threats, including integrity attacks and availability attacks.
@@
    Integrity attacks: The attacker’s goal is to tamper the integrity of the target. Specifically, this implies that the attack samples generated by the attacker are only effective in certain parts of the target system, while the remainder of the target system retains its original functionality.
@@
    Availability attacks: In contrast, availability attacks aim to disrupt the entire system, making it unavailable to legitimate users.
The main difference between integrity attacks and availability attacks lies in the focus of the optimization objectives of the attack models. In this paper, the proposed attack method is an availability attack, namely the destruction of the usability of all functions of a deep receiver.
@@ III-A2 Attacker’s Knowledge
The attacker’s knowledge indicates the extent to which they are aware of the attack target. This knowledge encompasses several dimensions, as outlined in [24],[36]. These dimensions include: (i) The data utilized for training purposes. (ii) The architectural design of the target model, the learning algorithms employed during training, and their associated parameters and hyperparameters. (iii) The data comprising the test set. Based on these dimensions, two main attack scenarios can be defined:
@@
    White-box attacks: The attacker has complete knowledge about the attack target. In this context, the attacker will adapt the nature of the attack to align with the specific characteristics of the target to achieve the most effective and impactful outcome.
@@
    Black-box attacks: Black-box attacks can be further categorized into two main types: Transfer-based attacks and query attacks. In transfer-based attacks, the attacker has limited or no knowledge of the target model. The knowledge about the target encompasses the aforementioned dimensions (i), (ii), and (iii). In this setting, the attacker is limited to relying on the data they have collected to construct a surrogate model that approximates the target model. This attack is then transferred to the target model by launching a white-box attack on the surrogate model. In a black-box query attacks, the attacker can query the target’s output or confidence level to optimize the attack. Currently, the majority of black-box attacks exploit the transferability for attack purposes [24],[38, 39, 40, 41].
@@ III-A3 Attacker’s Capability
The attacker’s capabilities determine the methods used to influence the attack target and the constraints for data manipulation. To avoid potential defense filtering mechanisms, the attacker typically imposes an upper bound of $\epsilon$ to the perturbation $\delta$ under the $p$-norm space. From the perspective of attack methods[24], if the attacker can manipulate data from both the training and testing phases, which considered as causal attacks and is called data poisoning attacks. If the attacker can only manipulate the data during the testing phase, this attack is considered exploratory and is called evasion attacks. The difference lies in the optimization objectives and implementation methods. This paper focuses on adversarial poisoning attacks in black-box scenarios, which can be seen as a synthesis of evasion attacks and data poisoning attacks. The specific optimization goals and implementation forms of evasion attacks and data poisoning attacks are described in the following Section III-B and Section III-C, respectively.
@@ III-B Evasion Attacks
The evasion attacks executed by generating adversarial samples aimed at causing misclassification during the testing phase of the DNN. This is achieved by identifying the network’s vulnerabilities and applying small, strategically crafted perturbations to the input, as illustrated in Fig. 5(a). Gradient-based optimizers [42],[43] are effective in determining these influential perturbations. Given a target model and input data, the gradient of the objective function is used to guide the application of minor perturbations, which maximizes the loss induced on the input. This can be formulated as a single-layer optimization problem. Specifically, let $\mathbf{y}$ represent the model input, $\mathbf{s}$ the corresponding labels, and $\mathbf{y}^{\prime}$ the adversarial samples generated after adding perturbation. The adversarial perturbation, denoted as $\delta=\left\|\mathbf{y}^{\prime}-\mathbf{y}\right\|_{p}$, is constrained by an upper bound $\epsilon>0$, with the optimal adversarial sample $\mathbf{y}^{*}$. The classifier is parameterized by $\theta$, and for classification tasks, the cross-entropy loss function $\mathcal{L}\left(\mathbf{y}^{\prime},\mathbf{s};f\theta\right)$ is employed. Thus, the generation of optimal adversarial samples can be framed as the following optimization problem